OpenStudy (anonymous):

Find a set of vectors {u, v} in R4 that spans the solution set of the equations: x−y+2z−3w=0 and x+2y+z+3w=0

OpenStudy (anonymous):

Really just want the answer at this point...please.

OpenStudy (anonymous):

I think you need to row reduce

OpenStudy (anonymous):

I tried row reducing, then parametric forms, etc it didnt get me anywhere, for me anyways.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/534202a7e4b09298524f1474 did you get that?

OpenStudy (anonymous):

The link seems to take me back to the question? :p

OpenStudy (anonymous):

O.O since when did copy this link? XD hold on ..

OpenStudy (anonymous):

Agh...no, I did not.

OpenStudy (anonymous):

ok, assume you did :D then x = (-5/3)z + w y = (1/3)z - 2w z = z w = w

OpenStudy (anonymous):

so the two vectors span of the two vectors are: z (-5/3, 1/3, 1,0) + w (1,-2,0,1)

OpenStudy (anonymous):

typo *the span of the two vectors*...

OpenStudy (anonymous):

Thank you for clearing that up. You should be called Sweetwing xD, But I understand now. Thanks again.

OpenStudy (anonymous):

ahaha sour wing sounds tastier XD. And you're welcome! ^.^